Calculate the Engery Needs for a woman using Harris Benedict who is 185#, 5’10”, and is 55 years old.

A male, 5’11” tall weighs 190# and is 29 years old.  Calculate his protein needs assuming he just had an appendectomy (minor surgery).
Answer #1

●Given:

Weight:185(83.9kg)

Height:5'10"(155cm)

Age:55years

Gender:Female

Harris Benedict

BMR=655.1+(4.35×Weight in pounds)+(1.850×height in cm)_4.676×age in years

=655.1+(4.35×185)+(1.850×155)_4.676×55

=655.1+(804.75)+(286.75)_257.18

=655.1+(804.75+286.75)_257.18

=655.1+(1091.5-257.18)

=655.1+834.32

=1489.42kcal/day

● Given

Gender:Male

Height:5'11"

Weight:190(86kg)

Age:29years

To calculate the amount of protein :multiply weight in pounds by 0.37

=190×0.37

=70.3gm of protein

After a surgery the protein requirements per kg for a minor surgery should be at least 1.2 to 2gm/kg

As the patient weighs 86kg ,multiply with the protein per kg after surgery

=86×1.2/ 86×2

=103.2 to 172 gm of protein per day to heal after a minor surgery.